Part A Calculate the equilibrium constant for the formation of NO2 (g) from NO (g) and O2 (g) at 405 K. 2 NO (g) + O2 (g) → 2 NO2 (g) Calculate the equilibrium constant for the formation of NO2 (g) from NO (g) and O2 (g) at 405 K. 2 NO (g) + O2 (g) → 2 NO2 (g)
Answer #1

    2 NO (g) + O2 (g) → 2 NO2 (g)

\DeltaH = Hp -HR

        = 2*33.2 -(2*90.2-0)   = -114Kj/mole

\DeltaS   = S0p - S0R

        = 2*240-(2*210.7+205) = -146.4J/mole-K   = -0.1464Kj/mole-K

\DeltaG0 = \DeltaH -T\DeltaS

          = -114-298*-0.1464   = -70.3728KJ/mole

\DeltaG0      = -RTlnK

-70372   = -8.314*405*298logK

logK    = -70372/-1003416.66

logK    = 0.07013

   K     = 100.07013   = 1.1752 >>>>>> answer